OpenStudy (anonymous):

Use a formula to find the surface area of the square pyramid. (1 point) 45 ft2 81 ft2 36 ft2 72 ft2

OpenStudy (anonymous):

if dimensions of the square are axa and it has height h then formula is $$ A = a^2 + 2\sqrt{\frac{a^2}{4} + h^2} $$